# Deep Code Search with Naming-Agnostic Contrastive Multi-View Learning

### v1

@@ Abstract.
Software development is a repetitive task, as developers usually reuse or get inspiration from existing implementations. Code search, which refers to the retrieval of relevant code snippets from a codebase according to the developer’s intent that has been expressed as a query, has become increasingly important in the software development process. Due to the success of deep learning in various applications, a great number of deep learning based code search approaches have sprung up and achieved promising results. However, developers may not follow the same naming conventions and the same variable may have different variable names in different implementations, bringing a challenge to deep learning based code search methods that rely on explicit variable correspondences to understand source code. To overcome this challenge, we propose a naming-agnostic code search method (NACS) based on contrastive multi-view code representation learning. NACS strips information bound to variable names from Abstract Syntax Tree (AST), the representation of the abstract syntactic structure of source code, and focuses on capturing intrinsic properties solely from AST structures. We use semantic-level and syntax-level augmentation techniques to prepare realistically rational data and adopt contrastive learning to design a graph-view modeling component in NACS to enhance the understanding of code snippets. We further model ASTs in a path view to strengthen the graph-view modeling component through multi-view learning. Extensive experiments show that NACS provides superior code search performance compared to baselines and NACS can be adapted to help existing code search methods overcome the impact of different naming conventions.
@@
ccs: Information systems Retrieval models and rankingccs: Software and its engineering Software notations and toolsccs: Computing methodologies Learning latent representations
@@ 1. Introduction
Code search takes search queries that manifest the software developer’ intent as inputs and returns desired code snippets. To enhance software development productivity and quality, programmers use code search tools to find high-quality code snippets in existing projects when they debug, write code, look for code to reuse, or learn API usage (Liu et al., 2022b). Existing studies show during software development, roughly one-fifth of the development time is used to search code examples (Brandt et al., 2009; Sun et al., 2022), indicating that code search has become an essential part of software development.
@@
Due to the pivotal role of code search in software development, much effort has been devoted to improving the quality of code search. Early works mainly use keyword matching between code snippets and search queries written in natural languages (Liu et al., 2022b). These methods typically employ traditional information retrieval algorithms (McMillan et al., 2011; Lv et al., 2015; Bajracharya et al., 2010) to measure the relevance between queries and candidate code snippets. Recently, the success of deep learning techniques has greatly promoted the development of code search approaches (Liu et al., 2022b; Yang et al., 2021; Zhang et al., 2023a). Various deep learning techniques have been introduced to improve code search, including but not limited to deep neural networks (Gu et al., 2018), meta learning (Chai et al., 2022), and pre-training (Feng et al., 2020; Guo et al., 2021). They encode queries and code snippets into low-dimension spaces and measure their relevance through the similarity (e.g., cosine similarity) between representation vectors.
@@
However, existing works lack the consideration of the impact of different naming styles (we call it naming issue in this paper for simplicity). Software developers may not follow the same naming convention, resulting in different naming of the same variable (Guo et al., 2021). The naming issue constitutes a barrier for effectively deploying deep learning techniques in code search since (1) it is hard to find correspondences when the same variable has different variable name in different code snippets, and (2) many deep learning techniques rely on entity correspondences (i.e., the same object is always represented by the same representation vector). In the literature, very few works have been done on remedying the naming issue. Guo et al. (2021) propose to model the data flow graph, which represents the dependency relation between variables and is the same under different abstract grammars for the same source code, to avoid the impact of different naming conventions in code representation learning. Nevertheless, source code is different from plain text as it requires strict grammar. How to avoid the impact of the naming issue when modeling the Abstract Syntax Tree (AST) [^1] The background of AST is provided in Section 3.1.1., which represents the abstract syntax of source code and is prevalently modeled in code representation learning (Allamanis et al., 2018a; Liu et al., 2022b; Yang et al., 2021), is under-exploited. Previous code representation learning methods not only capture syntactic information contained in ASTs but also encode AST node names, which contain variable names, as the semantic information of ASTs. Figure 1 provides an example to explain why AST-based code representation learning methods suffer from the naming issue. The left part of Figure 1 provides examples of two python code snippets. The code snippet at the bottom is the result after changing variable names in the code snippet on the top. The right part of Figure 1 shows ASTs of the two code snippets. The blue blocks correspond to the original variable names, and they are replaced by the green blocks after changing variable names. It is difficult for contemporary code representation learning techniques to understand that corresponding variables in the two code snippets (e.g., lst and l) are the same due to their different names (Guo et al., 2021).
Figure 1. Examples of two code snippets and their corresponding ASTs.
@@ 2.1. Code Representation Learning
Software is ubiquitous in modern society and there is an ongoing demand for developing new software and enhancing existing software (Yang et al., 2021). Considering the importance of software development, a great number of assistance tools (e.g., code completion tool (Svyatkovskiy et al., 2019; Wang and Li, 2021), code search tool (Lv et al., 2015; Zhang et al., 2023a), code summarization tool (Fowkes et al., 2016; Lin et al., 2021b), issue-commit recovery tool (Zhang et al., 2023b) and code refactoring tool (Liu et al., 2023)) have been developed to boost the efficiency of software development. Code and its affiliated documents (e.g., code comments) are important elements in software development. Strengthening the understanding of code can improve assistance tools and help engineers construct better software (Le et al., 2020). Inspired by the success of deep learning techniques, researchers have recently explored applying deep neural networks (DNNs) to enhance code representation learning and these works can be roughly classified into two categories.
@@ 2.2. Code Search
Code search, which is also called code retrieval in the literature, refers to the retrieval of relevant code snippets from a large codebase according to programmer’s intent that has been expressed as a query (Liu et al., 2022b; Xie et al., 2024). Code search lowers the cost of learning the usage of new APIs by providing code examples. Code search also reduces programmers’ cognitive burden and boosts the efficiency of software development by providing reusable code snippets that can be easily adapted. Therefore, code search has become increasing important in the software development process and attracted considerable attention.
@@
Graph supervised/semi-supervised learning relies on scarce data labels to supervise model learning. Differently, graph self-supervised learning (GSSL) (Wu et al., 2022) models graphs through handcrafted auxiliary tasks (pretext tasks) that acquire supervision signals from intrinsic graph properties (Liu et al., 2022a). Since GSSL alleviates label reliance, it draws more and more attention recently (Xie et al., 2022; Wu et al., 2021).
@@
  • Auxiliary property based methods adopt traditional graph tasks such as clustering (You et al., 2020b; Sun et al., 2020a), graph partitioning (You et al., 2020b), shortest distance prediction (Peng et al., 2020) or node similarity prediction (Zhu et al., 2020b; Jin et al., 2021) as the pretext task. These methods have a similar training paradigm with supervised learning since both of them learn with “sample-label” pairs (Liu et al., 2022a). The difference is auxiliary property based graph SSL generates pseudo label without manual labeling.
  • Contrast based methods are based on mutual information (MI) maximization (Hjelm et al., 2019). They maximize the MI between related graph instances and minimize the MI between unrelated graph instances (Hjelm et al., 2019; Liu et al., 2022a). Two stages are typically required for contrast based graph SSL. In the graph augmentation stage, different graph instances are generated through node feature masking (Jin et al., 2020), node feature shuffling (Ren et al., 2019), edge dropping/adding (You et al., 2020a) or subgraph sampling (Hassani and Ahmadi, 2020; Hu et al., 2020; Jiao et al., 2020; Qiu et al., 2020). In the graph contrastive learning stage, pretext tasks are leveraged to maximize the MI between positive instances. Contrast can be conducted at same scale or cross scales. The former discriminates instances in an equal scale (e.g., node versus node) (Qiu et al., 2020; You et al., 2020a) while the later discriminates instances across different granularities (e.g., node versus subgraph) (Mavromatis and Karypis, 2021; Robinson et al., 2021).
  • Compared to methods that only consider one pretext task, hybrid methods adopt several pretext tasks to leverage different types of supervisions (Liu et al., 2022a; You et al., 2020b). These methods optimize various pretext tasks in a manner of multi-task learning.
@@ 3.1.1. Abstract Syntax Tree
An Abstract Syntax Tree (AST) of a code snippet $c$ is a tuple $\mathcal{NT},\mathcal{T},r,f\right>$ where $\mathcal{NT}$ is the non-terminal node set, $\mathcal{T}$ is the terminal node set, $r$ is the root node, and $f:\mathcal{NT}\rightarrow(\mathcal{NT}\cup\mathcal{T})$ is a mapping function that maps a non-terminal node to its child nodes.
@@ 3.1.2. PT and ST strategies
ST Strategies. In addition to modeling the semantic meaning of code snippets, capturing the structural features by learning the AST of each code snippet as a graph is commonly adopted in code representation learning (Wang and Li, 2021; Allamanis et al., 2018b). As ASTs represent syntactic meaning of code snippets, structure-level learning is equivalent to capturing syntactic information. Recently, graph contrastive learning has attracted considerable attention and achieves promising performance in various graph based applications (Wu et al., 2021). Nevertheless, existing graph data augmentation strategies are not suitable for code representation learning. We design two graph transformation methods to generate structure-augmented data of the AST of each code snippet, which enhances the structure-level understanding of NACS:
@@ 3.2.1. Graph-View Modeling
[…]
@@
    where $\operatorname{MLP}(\cdot)^{(k)}$ denotes the multi-layer percetron for the $k$-$th$ layer, $\mathbf{h}_{s}^{(k)}$ represents the output for the AST node $s$ at the $k$-$th$ layer of GIN, $\mathcal{N}(s)$ is the set of neighbors of $s$, $\epsilon$ is a learnable parameter, and we set $\mathbf{h}_{s}^{(0)}=\mathbf{b}_{s}$. Then, we perform a mean pooling operation on representations of all nodes in $g$ output by GIN. The result is fed to a MLP to get the graph representation $\mathbf{r}_{g}^{(i)}$ of $g$ from the $i$-$th$ layer. The final graph representation $\textbf{r}_{g}$ of the AST graph $g$ is the sum of $\{\mathbf{r}_{g}^{(1)}\cdots\mathbf{r}_{g}^{(K)}\}$ where $K$ is the number of layers:
@@ 3.2.2. Path-View Modeling
For each AST $x_{c}^{q}$ of a code snippet $c$, we extract all its paths from the root node to the leaf node =$\mathcal{R}=\{y_{c,1},y_{c,2},\cdots,y_{c,z_{c}}\}$ where $z_{c}$ represents the total number of paths in $x_{c}^{q}$. Since the numbers of paths in different ASTs differ, $z_{c}$ is not fixed. $y_{c,i}=\{n_{c,i,1},n_{c,i,2},\cdots,n_{c,i,l_{i}}\}$, where $l_{i}$ represents the path length of the AST path $i$ and $n_{c,i,j}$ ($1\leq j\leq l_{i}$) is the AST node type of the $j$-th node on the AST path $i$. We apply a Bi-LSTM, which can extract information in forward and reverse directions, to encode the AST path. For an AST path $i$ in $x_{c}^{q}$, at each time step $s$, the Bi-LSTM reads the embedding of the $s$-th node in $i$, then computes the hidden states $\mathbf{h}_{c,i,s}$:
@@
Taking one positive sample $x_{c}^{p1}$ out of the three positive samples ($x_{c}^{p1}$, $x_{c}^{p2}$ and $x_{c}^{p3}$) for an AST $x_{c}^{q}$ as an example, in path-view modeling, the contrastive loss of the positive pair $x_{c}^{q},\,x_{c}^{p1}\right>$ is defined as follows:
@@ 4.1.1. Data
We choose three public datasets that are commonly used in previous works in our experiments:
@@
    CodeSearchNet dataset [^2] https://github.com/github/CodeSearchNet (Husain et al., 2019): It is the Python dataset used in the CodeSearchNet challenge. It contains 0.5M bimodal Python functions and their corresponding natural language descriptions, as well as 1.1M unimodal python functions. CodeSearchNet dataset is widely used in various code relevant tasks including but not limited to code representation learning (Feng et al., 2020; Guo et al., 2021), code summarization (Lin et al., 2021b) and code completion (Ciniselli et al., 2021).
@@
We extract ASTs using Python official AST library [^4] https://docs.python.org/3/library/ast.html.
@@ 4.1.3. Baselines
4.1.4. Hyper-Parameter Setting and Experiment Environment

### v2

@@ Abstract.
Software development is a repetitive task, as developers usually reuse or get inspiration from existing implementations. Code search, which refers to the retrieval of relevant code snippets from a codebase according to the developer’s intent that has been expressed as a query, has become increasingly important in the software development process. Due to the success of deep learning in various applications, a great number of deep learning based code search approaches have sprung up and achieved promising results. However, developers may not follow the same naming conventions and the same variable may have different variable names in different implementations, bringing a challenge to deep learning based code search methods that rely on explicit variable correspondences to understand source code. To overcome this challenge, we propose a naming-agnostic code search method (NACS) based on contrastive multi-view code representation learning. NACS strips information bound to variable names from Abstract Syntax Tree (AST), the representation of the abstract syntactic structure of source code, and focuses on capturing intrinsic properties solely from AST structures. We use semantic-level and syntax-level augmentation techniques to prepare realistically rational data and adopt contrastive learning to design a graph-view modeling component in NACS to enhance the understanding of code snippets. We further model ASTs in a path view to strengthen the graph-view modeling component through multi-view learning. Extensive experiments show that NACS provides superior code search performance compared to baselines and NACS can be adapted to help existing code search methods overcome the impact of different naming conventions. Our implementation is available at https://github.com/KDEGroup/NACS.
@@
$c$s: Information systems Retrieval models and rankingccs: Software and its engineering Software notations and toolsccs: Computing methodologies Learning latent representations
@@ 1. Introduction
Code search takes search queries that manifest the software developer’ intent as inputs and returns desired code snippets. To enhance software development productivity and quality, programmers use code search tools to find high-quality code snippets in existing projects when they debug, write code, look for code to reuse, or learn API usage (Liu et al., 2022). Existing studies show during software development, roughly one-fifth of the development time is used to search code examples (Brandt et al., 2009; Sun et al., 2022), indicating that code search has become an essential part of software development.
@@
Due to the pivotal role of code search in software development, much effort has been devoted to improving the quality of code search. Early works mainly use keyword matching between code snippets and search queries written in natural languages (Liu et al., 2022). These methods typically employ traditional information retrieval algorithms (McMillan et al., 2011; Lv et al., 2015; Bajracharya et al., 2010) to measure the relevance between queries and candidate code snippets. Recently, the success of deep learning techniques has greatly promoted the development of code search approaches (Liu et al., 2022; Yang et al., 2022; Zhang et al., 2023a). Various deep learning techniques have been introduced to improve code search, including but not limited to deep neural networks (Gu et al., 2018), meta learning (Chai et al., 2022), and pre-training (Feng et al., 2020; Guo et al., 2021). They encode queries and code snippets into low-dimension spaces and measure their relevance through the similarity (e.g., cosine similarity) between representation vectors.
@@
However, existing works lack the consideration of the impact of different naming styles (we call it naming issue in this paper for simplicity). Software developers may not follow the same naming convention, resulting in different naming of the same variable (Guo et al., 2021). The naming issue constitutes a barrier for effectively deploying deep learning techniques in code search since (1) it is hard to find correspondences when the same variable has different variable name in different code snippets, and (2) many deep learning techniques rely on entity correspondences (i.e., the same object is always represented by the same representation vector). In the literature, very few works have been done on remedying the naming issue. Guo et al. (2021) propose to model the data flow graph, which represents the dependency relation between variables and is the same under different abstract grammars for the same source code, to avoid the impact of different naming conventions in code representation learning. Nevertheless, source code is different from plain text as it requires strict grammar. How to avoid the impact of the naming issue when modeling the Abstract Syntax Tree (AST) [^1] The background of AST is provided in Section 3.1.1., which represents the abstract syntax of source code and is prevalently modeled in code representation learning (Allamanis et al., 2018a; Liu et al., 2022; Yang et al., 2022), is under-exploited. Previous code representation learning methods not only capture syntactic information contained in ASTs but also encode AST node names, which contain variable names, as the semantic information of ASTs. Figure 1 provides an example to explain why AST-based code representation learning methods suffer from the naming issue. The left part of Figure 1 provides examples of two python code snippets. The code snippet at the bottom is the result after changing variable names in the code snippet on the top. The right part of Figure 1 shows ASTs of the two code snippets. The blue blocks correspond to the original variable names, and they are replaced by the green blocks after changing variable names. It is difficult for contemporary code representation learning techniques to understand that corresponding variables in the two code snippets (e.g., lst and l) are the same due to their different names (Guo et al., 2021).
@@
Figure 1. Examples of two code snippets and their corresponding ASTs.
@@ 2.1. Code Representation Learning
Software is ubiquitous in modern society and there is an ongoing demand for developing new software and enhancing existing software (Yang et al., 2022). Considering the importance of software development, a great number of assistance tools (e.g., code completion tool (Svyatkovskiy et al., 2019; Wang and Li, 2021), code search tool (Lv et al., 2015; Zhang et al., 2023a), code summarization tool (Fowkes et al., 2016; Lin et al., 2021b), issue-commit recovery tool (Zhang et al., 2023b) and code refactoring tool (Liu et al., 2023b)) have been developed to boost the efficiency of software development. Code and its affiliated documents (e.g., code comments) are important elements in software development. Strengthening the understanding of code can improve assistance tools and help engineers construct better software (Le et al., 2020). Inspired by the success of deep learning techniques, researchers have recently explored applying deep neural networks (DNNs) to enhance code representation learning and these works can be roughly classified into two categories.
@@ 2.2. Code Search
Code search, which is also called code retrieval in the literature, refers to the retrieval of relevant code snippets from a large codebase according to programmer’s intent that has been expressed as a query (Liu et al., 2022; Xie et al., 2024). Code search lowers the cost of learning the usage of new APIs by providing code examples. Code search also reduces programmers’ cognitive burden and boosts the efficiency of software development by providing reusable code snippets that can be easily adapted. Therefore, code search has become increasing important in the software development process and attracted considerable attention.
@@
Graph supervised/semi-supervised learning relies on scarce data labels to supervise model learning. Differently, graph self-supervised learning (GSSL) (Wu et al., 2022) models graphs through handcrafted auxiliary tasks (pretext tasks) that acquire supervision signals from intrinsic graph properties (Liu et al., 2023a). Since GSSL alleviates label reliance, it draws more and more attention recently (Xie et al., 2023; Wu et al., 2023).
@@
  • Auxiliary property based methods adopt traditional graph tasks such as clustering (You et al., 2020b; Sun et al., 2020a), graph partitioning (You et al., 2020b), shortest distance prediction (Peng et al., 2020) or node similarity prediction (Zhu et al., 2020b; Jin et al., 2021) as the pretext task. These methods have a similar training paradigm with supervised learning since both of them learn with “sample-label” pairs (Liu et al., 2023a). The difference is auxiliary property based graph SSL generates pseudo label without manual labeling.
  • Contrast based methods are based on mutual information (MI) maximization (Hjelm et al., 2019). They maximize the MI between related graph instances and minimize the MI between unrelated graph instances (Hjelm et al., 2019; Liu et al., 2023a). Two stages are typically required for contrast based graph SSL. In the graph augmentation stage, different graph instances are generated through node feature masking (Jin et al., 2020), node feature shuffling (Ren et al., 2019), edge dropping/adding (You et al., 2020a) or subgraph sampling (Hassani and Ahmadi, 2020; Hu et al., 2020; Jiao et al., 2020; Qiu et al., 2020). In the graph contrastive learning stage, pretext tasks are leveraged to maximize the MI between positive instances. Contrast can be conducted at same scale or cross scales. The former discriminates instances in an equal scale (e.g., node versus node) (Qiu et al., 2020; You et al., 2020a) while the later discriminates instances across different granularities (e.g., node versus subgraph) (Mavromatis and Karypis, 2021; Robinson et al., 2021).
  • Compared to methods that only consider one pretext task, hybrid methods adopt several pretext tasks to leverage different types of supervisions (Liu et al., 2023a; You et al., 2020b). These methods optimize various pretext tasks in a manner of multi-task learning.
@@ 3.1.1. Abstract Syntax Tree
An Abstract Syntax Tree (AST) of a code snippet $c$ is a tuple $\left<\mathcal{NT},\mathcal{T},r,f\right$> where $\mathcal{NT}$ is the non-terminal node set, $\mathcal{T}$ is the terminal node set, $r$ is the root node, and $f:\mathcal{NT}\rightarrow(\mathcal{NT}\cup\mathcal{T})$ is a mapping function that maps a non-terminal node to its child nodes.
@@ 3.1.2. PT and ST Strategies
ST Strategies. In addition to modeling the semantic meaning of code snippets, capturing the structural features by learning the AST of each code snippet as a graph is commonly adopted in code representation learning (Wang and Li, 2021; Allamanis et al., 2018b). As ASTs represent syntactic meaning of code snippets, structure-level learning is equivalent to capturing syntactic information. Recently, graph contrastive learning has attracted considerable attention and achieves promising performance in various graph based applications (Wu et al., 2023). Nevertheless, existing graph data augmentation strategies are not suitable for code representation learning. We design two graph transformation methods to generate structure-augmented data of the AST of each code snippet, which enhances the structure-level understanding of NACS:
@@ 3.2.1. Graph-View Modeling
    For each node $s$ in an AST graph $g$, we encode its AST node type $\mathbf{b}_{s}^{\text{type}}$ and its degree $\mathbf{b}_{s}^{\text{degree}}$ (i.e., all dimensions of $\mathbf{b}_{s}^{\text{degree}}$ are the degree of $s$ in the AST graph) as part of node features. Additionally, following the initialization method used in existing GNN pre-training approaches (Qiu et al., 2020), we conduct eigen-decomposition on each AST graph’s normalized graph Laplacian,  $\textit{ s.t. }\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top}$, and then add the top eigenvectors of $\mathbf{U}$ to construct node features of $s$:
@@
    where $\operatorname{MLP}(\cdot)^{(k)}$ denotes the multi-layer percetron for the $k$-thth layer, $\mathbf{h}_{s}^{(k)}$ represents the output for the AST node $s$ at the $k$-thth layer of GIN, $\mathcal{N}(s)$ is the set of neighbors of $s$, $\epsilon$ is a learnable parameter, and we set $\mathbf{h}_{s}^{(0)}=\mathbf{b}_{s}$. Then, we perform a mean pooling operation on representations of all nodes in $g$ output by GIN. The result is fed to a MLP to get the graph representation $\mathbf{r}_{g}^{(i)}$ of $g$ from the $i$-thth layer. The final graph representation $\textbf{r}_{g}$ of the AST graph $g$ is the sum of $\{\mathbf{r}_{g}^{(1)}\cdots\mathbf{r}_{g}^{(K)}\}$ where $K$ is the number of layers:
@@ 3.2.2. Path-View Modeling
For each AST $x_{c}^{q}$ of a code snippet $c$, we extract all its paths from the root node to the leaf node $\mathcal{R}=\{y_{c,1},y_{c,2},\cdots,y_{c,z_{c}}\}$ where $z_{c}$ represents the total number of paths in $x_{c}^{q}$. Since the numbers of paths in different ASTs differ, $z_{c}$ is not fixed. $y_{c,i}=\{n_{c,i,1},n_{c,i,2},\cdots,n_{c,i,l_{i}}\}$, where $l_{i}$ represents the path length of the AST path $i$ and $n_{c,i,j}$ ($1\leq j\leq l_{i}$) is the AST node type of the $j$-th node on the AST path $i$. We apply a Bi-LSTM, which can extract information in forward and reverse directions, to encode the AST path. For an AST path $i$ in $x_{c}^{q}$, at each time step $s$, the Bi-LSTM reads the embedding of the $s$-th node in $i$, then computes the hidden states $\mathbf{h}_{c,i,s}$:
@@
Taking one positive sample $x_{c}^{p1}$ out of the three positive samples ($x_{c}^{p1}$, $x_{c}^{p2}$ and $x_{c}^{p3}$) for an AST $x_{c}^{q}$ as an example, in path-view modeling, the contrastive loss of the positive pair $\left<x_{c}^{q},\,x_{c}^{p1}\right$> is defined as follows:
@@ 4.1.1. Data
We choose four public datasets that are commonly used in previous works in our experiments:
@@
    CodeSearchNet-Python dataset [^2] https://github.com/github/CodeSearchNet (Husain et al., 2019): It is the Python dataset used in the CodeSearchNet challenge. It contains 0.5M Python functions and their corresponding natural language descriptions. CodeSearchNet dataset is widely used in various code relevant tasks including but not limited to code representation learning (Feng et al., 2020; Guo et al., 2021), code summarization (Lin et al., 2021b) and code completion (Ciniselli et al., 2021).
  2. (2)
    CodeSearchNet-Java dataset [^3] https://huggingface.co/datasets/Nan-Do/code-search-net-java (Husain et al., 2019): It is the Java dataset used in the CodeSearchNet challenge. It contains 0.5M Java functions and their corresponding natural language descriptions.
@@ 4.1.3. Baselines
4.1.4. Implementation Details
For Python datasets CodeSearchNet-Java, CoSQA and CoSQA-Var, we extract ASTs and conduct data augmentation for NACS using Python official AST library [^11] https://docs.python.org/3/library/ast.html. For Java dataset CodeSearchNet-Java, we extract ASTs and conduct data augmentation for NACS using JavaParser library [^12] https://javaparser.org. We use the implementations from authors for baselines. We implement NACS using PyTorch. The implementation and datasets are available at https://github.com/KDEGroup/NACS.